OpenStudy (anonymous):

Is there ANY number, when it is substituted in the quadratic formula (x2 - x + 41), the result is a composite number? Please explain how to find this number or the rule that makes no number possible...THANK YOU!

hero (hero):

What?

OpenStudy (anonymous):

Here's the assignment: "Mathematicians have been searching for a formula that yields prime numbers. One such formula was x2 - x + 41. Select some numbers for x, substitute them in the formula, and see if prime numbers occur. Try to find a number for x that when substituted in the formula yields a composite number." I am trying to do the second part-substituting a number for x that yields a composite number.

OpenStudy (anonymous):

hmm it occurs to me that you might need to search a number so that one can factorize something from the formula, 41 might do the job

OpenStudy (anonymous):

so 〖 41〗^2-41+41=41×41=1681...what are the factor of 1681?

OpenStudy (anonymous):

oh wait, 41 x 41

OpenStudy (anonymous):

lol